To resolve an object in an electron microscope, the wavelength of the electrons must be close to the diameter of the object. What kinetic energy, ?k , must the electrons have in order to resolve a protein molecule that is 5.90 nm in diameter? Take the mass of an electron to be 9.11×10−31 kg

?k= ______J
Answer #1

Here from the de Broglie wavelength we have the following relation as,

Lambda = h/mv

Where

h = Planck's constant = 6.626 x 10^-34 Js

m = 9.11x 10-31 kg

Lambda = 5.9 x 10^-9 m

Also here Electron velocity will be given as,

v = h/m×lambda

V = (6.626 x 10^-34)/(9.11x 10^–31 )(5.9x10^-9)

V = (123276.71212487674) m/s

Now here kinetic energy will be given as,

Ek = (1/2) mv^2 = (1/2) (9.11× 10-31 kg) (123276.712)^2

E = 6922.300 x 10^-24 J

Hence,

Ek = 6922.300 x 10^-24 J